Please find the equation with a x-intercept of 5 and a slope of 3.
Answer by nerdybill(7384)

Please find the equation with a x-intercept of 5 and a slope of 3.
.
Slope-intercept form of a line:
y = mx + b
.
The problem gives you:
m = 5
and x-intercept of 5 or (x,y) = (5,0)
.
Plug everything in to find 'b':
0 = 5(5) + b
0 = 25 + b
-25 = b
.
Therefore, your equation is
y = 5x - 25